# Q-SINDy: Quantum-Kernel Sparse Identification of Nonlinear Dynamics with Provable Coefficient Debiasing

### 21 Apr 2026 v2

@@ Abstract
Quantum feature maps offer expressive embeddings for classical learning tasks, and augmenting sparse identification of nonlinear dynamics (SINDy) with such features is a natural but unexplored direction. We introduce Q-SINDy, a quantum-kernel-augmented SINDy framework, and identify a specific failure mode that arises: coefficient cannibalization, in which quantum features absorb coefficient mass that rightfully belongs to the polynomial basis, corrupting equation recovery. We derive the exact cannibalization-bias formula $\Delta\xi_{P}=(P^{\top}P)^{-1}P^{\top}Q\,\hat{\xi}_{Q}$ and prove that orthogonalizing quantum features against the polynomial column space at fit time eliminates this bias exactly. The claim is verified numerically to machine precision ($10^{-12}$) on multiple systems. Empirically, across six canonical dynamical systems (Duffing, Van der Pol, Lorenz, Lotka-Volterra, cubic oscillator, Rössler) and three quantum feature map architectures (ZZ-angle encoding, IQP, data re-uploading), orthogonalized Q-SINDy consistently matches vanilla SINDy’s structural recovery while uncorrected augmentation degrades true-positive rates by up to 100%. A refined dynamics-aware diagnostic, $R^{2}_{Q}$ for $\dot{X}$, predicts cannibalization severity with statistical significance (Pearson $r=0.70$, $p=0.023$). An RBF classical-kernel control across 20 hyperparameter configurations fails more severely than any quantum variant, ruling out feature count as the cause. Orthogonalization remains robust under depolarizing hardware noise up to 2% per gate, and the framework extends without modification to Burgers’ equation.
@@ 1 Introduction
    Empirical validation spans six dynamical systems, three quantum feature map architectures (ZZ-angle, IQP, data re-uploading), hardware-noise simulation with depolarizing channels, and a PDE extension to Burgers’ equation.
@@ 3.2 Quantum feature maps
ZZ angle-encoded (2- or 3-qubit). Data-encoding rotations $R_{X}(x_{i})$ followed by $R_{ZZ}(x_{i}x_{j})$ pairwise entanglers and a second rotation layer $R_{Y}(x_{i})$, measured in Pauli-$Z$, Pauli-$X$, and $ZZ/XX$ bases.
@@ 5.5 Generality across quantum feature-map architectures
We test whether the cannibalization+orthogonalization pattern generalizes beyond the ZZ-angle encoding. On Duffing and Van der Pol, we swap in IQP and data re-uploading feature maps. Table 3 reports $R^{2}_{Q}$ and observed cannibalization severity for each (system, feature-map) combination; orthogonalized Q-SINDy matches vanilla across all six ZZ/IQP/re-up variants on these two systems, with residual polynomial-coefficient deviations $\mathcal{O}(10^{-14})$ in every case. The mechanism is not specific to any particular quantum architecture. Circuit specifications for each feature map appear in Appendix B.
@@ 6 Discussion and Limitations
(1) All quantum feature maps tested (ZZ, IQP with 2 layers, data re-uploading) are classically simulable at the qubit counts we use; we make no claim of quantum computational advantage. Our contribution is methodological. (2) The $R^{2}_{Q}$ diagnostic was validated on $N=10$ $(\text{system},\text{feature-map})$ points; while leave-$k$-out cross-validation (Table 1) indicates the relationship generalizes within this set, scaling to 30+ combinations would tighten the statistical claim. (3) We demonstrate framework extension on one PDE (Burgers); cannibalization did not occur there because $R^{2}_{Q}$ was low. A PDE where cannibalization manifests—reaction–diffusion and Kuramoto–Sivashinsky are candidates—remains future work. (4) All experiments use STLSQ; whether cannibalization manifests equivalently under LASSO, Weak-SINDy, or Bayesian SINDy is unverified, though Proposition 4’s block-decoupling argument suggests orthogonalization should remain beneficial under any sparse regression that admits the same active-set normal equations.
@@ Appendix B Feature-map details
ZZ angle-encoded (2-qubit).
@@
ZZ angle-encoded (3-qubit).
Analogous 3-qubit generalization with pairwise ZZ entanglers in a ring, yielding 9 observables. Input data is rescaled by $\pi/(2\cdot\max|X|)$ for the 3D systems (Lorenz, Rössler, Burgers).
@@
Two layers of $H^{\otimes 2}$, $R_{Z}(x_{0}),R_{Z}(x_{1})$, $R_{ZZ}(x_{0}x_{1})$. Measurements as for ZZ.

### 22 Apr 2026 v3

@@ Abstract
Quantum feature maps offer expressive embeddings for classical learning tasks, and augmenting sparse identification of nonlinear dynamics (SINDy) with such features is a natural but unexplored direction. We introduce Q-SINDy, a quantum-kernel-augmented SINDy framework, and identify a specific failure mode that arises: coefficient cannibalization, in which quantum features absorb coefficient mass that rightfully belongs to the polynomial basis, corrupting equation recovery. We derive the exact cannibalization-bias formula $\Delta\xi_{P}=(P^{\top}P)^{-1}P^{\top}Q\,\hat{\xi}_{Q}$ and prove that orthogonalizing quantum features against the polynomial column space at fit time eliminates this bias exactly. The claim is verified numerically to machine precision ($10^{-12}$) on multiple systems. Empirically, across six canonical dynamical systems (Duffing, Van der Pol, Lorenz, Lotka-Volterra, cubic oscillator, Rössler) and three quantum feature map architectures ($Z$-angle encoding, IQP, data re-uploading), orthogonalized Q-SINDy consistently matches vanilla SINDy’s structural recovery while uncorrected augmentation degrades true-positive rates by up to 100%. A refined dynamics-aware diagnostic, $R^{2}_{Q}$ for $\dot{X}$, predicts cannibalization severity with statistical significance (Pearson $r=0.70$, $p=0.023$). An RBF classical-kernel control across 20 hyperparameter configurations fails more severely than any quantum variant, ruling out feature count as the cause. Orthogonalization remains robust under depolarizing hardware noise up to 2% per gate, and the framework extends without modification to Burgers’ equation.
@@ 1 Introduction
    Empirical validation spans six dynamical systems, three quantum feature map architectures ($Z$-angle, IQP, data re-uploading), hardware-noise simulation with depolarizing channels, and a PDE extension to Burgers’ equation.
@@ 3.2 Quantum feature maps
$Z$ angle-encoded (2- or 3-qubit). Data-encoding rotations $R_{X}(x_{i})$ followed by $R_{ZZ}(x_{i}x_{j})$ pairwise entanglers and a second rotation layer $R_{Y}(x_{i})$, measured in Pauli-$Z$, Pauli-$X$, and $ZZ/XX$ bases.
@@ 5.5 Generality across quantum feature-map architectures
We test whether the cannibalization+orthogonalization pattern generalizes beyond the $Z$-angle encoding. On Duffing and Van der Pol, we swap in IQP and data re-uploading feature maps. Table 3 reports $R^{2}_{Q}$ and observed cannibalization severity for each (system, feature-map) combination; orthogonalized Q-SINDy matches vanilla across all six ZZ/IQP/re-up variants on these two systems, with residual polynomial-coefficient deviations $\mathcal{O}(10^{-14})$ in every case. The mechanism is not specific to any particular quantum architecture. Circuit specifications for each feature map appear in Appendix B.
@@ 6 Discussion and Limitations
(1) All quantum feature maps tested ($Z$, IQP with 2 layers, data re-uploading) are classically simulable at the qubit counts we use; we make no claim of quantum computational advantage. Our contribution is methodological. (2) The $R^{2}_{Q}$ diagnostic was validated on $N=10$ $(\text{system},\text{feature-map})$ points; while leave-$k$-out cross-validation (Table 1) indicates the relationship generalizes within this set, scaling to 30+ combinations would tighten the statistical claim. (3) We demonstrate framework extension on one PDE (Burgers); cannibalization did not occur there because $R^{2}_{Q}$ was low. A PDE where cannibalization manifests—reaction–diffusion and Kuramoto–Sivashinsky are candidates—remains future work. (4) All experiments use STLSQ; whether cannibalization manifests equivalently under LASSO, Weak-SINDy, or Bayesian SINDy is unverified, though Proposition 4’s block-decoupling argument suggests orthogonalization should remain beneficial under any sparse regression that admits the same active-set normal equations.
@@ Appendix B Feature-map details
$Z$ angle-encoded (2-qubit).
@@
$Z$ angle-encoded (3-qubit).
Analogous 3-qubit generalization with pairwise $Z$ entanglers in a ring, yielding 9 observables. Input data is rescaled by $\pi/(2\cdot\max|X|)$ for the 3D systems (Lorenz, Rössler, Burgers).
@@
Two layers of $H^{\otimes 2}$, $R_{Z}(x_{0}),R_{Z}(x_{1})$, $R_{ZZ}(x_{0}x_{1})$. Measurements as for $Z$.